### v2

@@ Appendix H Centaur Experiments
For humans, accuracy nudges did not improve truth discernment when sharing entertaining news (Figure 19, Panel A). As in [43], there were no significant differences in sharing intentions between the Control and Active Control conditions (Headline Veracity $\times$ Intervention: $b=0.14$, 95% CI [0.07, 0.34][-0.07,\,0.34], $t(2997)=1.31$, $p=.56$). In contrast to the previous findings, neither accuracy nudge significantly increased sharing discernment relative to the controls (e.g., Headline Veracity $\times$ Intervention: Treatment vs. Control, b=0.05b=-0.05, 95% CI [0.25, 0.16][-0.25,\,0.16], t(2997)=0.44t(2997)=-0.44, $p=.97$; Importance Treatment vs. Control, b=0.03b=-0.03, 95% CI [0.23, 0.17][-0.23,\,0.17], t(2997)=0.30t(2997)=-0.30, $p=.99$). See Table LABEL:SI_accuracy_tab3 for the full regression results.
@@
For digital twins, however, accuracy nudges successfully improved truth discernment when sharing entertaining news (Figure 19, Panel B). As in [43], there were no significant differences in sharing intentions between the Control and Active Control conditions (Headline Veracity $\times$ Intervention: b=0.06b=-0.06, 95% CI [0.26, 0.14][-0.26,\,0.14], t(2997)=0.58t(2997)=-0.58, $p=.94$), and both treatments significantly increased sharing discernment relative to the controls (e.g., Headline Veracity $\times$ Intervention: Treatment vs. Control, b=0.61b=-0.61, 95% CI [0.82,0.41][-0.82,\,-0.41], t(2997)=5.93t(2997)=-5.93, $p<.001$; Importance Treatment vs. Control, b=0.44b=-0.44, 95% CI [0.64,0.24][-0.64,\,-0.24], t(2997)=4.27t(2997)=-4.27, $p<.001$). See Table LABEL:SI_accuracy_tab4 for the full regression results.
@@
Table 7: Model B: Headline veracity, intervention, and participant type on sharing. * p<<0.05, ** p<<0.01, *** p<<0.001.
@@
As an additional test, [43] analysis plan was next replicated for humans and their digital twins, separately. While the pre-registered analysis demonstrates sharing intentions for the full range of responses, [43] rescale the continuous sharing intentions into a binary variable capturing responses at or above the midpoint at “likely to share” (= 1) and below the midpoint as “unlikely to share” (= 0), and filtered out participants who would not consider sharing these articles on social media. Because the human dataset revealed significant differences between the Control and Active Control conditions, these groups were not pooled to maintain comparability. This results in a slight deviation from the analysis in [43]. Using this framework, their exact analysis was replicated for humans and their digital twins separately, treating the rescaled likelihood to share variable as the DV with Headline Veracity condition, Intervention condition, and their interaction as predictors, applying robust standard errors. Figure 20 shows proportion of sharing across all conditions for humans (Panel A) and their digital twins (Panel B).
@@
Figure 20: Proportion (%) Likely to Share by Headline Veracity and Intervention. NOTE: Shown here is the fraction of “likely’ responses (responses above the midpoint of the six-point Likert scale) by Headline Veracity and Intervention condition; the full distributions of responses are shown below, in figure 21. As with [43], these analyses focus only on participants who indicated that they would consider sharing this social media content. Analysis including all participants does not change the results. Error bars indicate 95% confidence intervals based on robust standard errors clustered on participant and headline.
@@
Replicating the main findings for humans (Model A), accuracy nudges did not improve truth discernment when sharing entertaining news (Figure 20, Panel A). In contrast to [43], sharing discernment was marginally higher in the Active Control compared to the Control condition (Headline Veracity $\times$ Intervention: b=0.03b=-0.03, 95% CI [0.07, 0.00][-0.07,\,0.00], $F(1,5132)=3.53$, $p=.06$), and neither treatment significantly increased sharing discernment relative to the controls (Headline Veracity $\times$ Intervention: Treatment, $b=0.01$, 95% CI [0.05, 0.06][-0.05,\,0.06], $F(1,9656)=0.05$, $p=.82$; Importance Treatment, $b=0.01$, 95% CI [0.04, 0.05][-0.04,\,0.05], $F(1,9656)=0.06$, $p=.80$). See Table 21 for the full regression results.
Figure 21: Headline Veracity, Intervention and Participant Type on Humans’ Sharing Intentions.
@@
Again, replicating the main findings for digital twins (Model A), accuracy nudges improved truth discernment when sharing entertaining news (Figure 20, Panel B). As in [43], there were no significant differences in sharing intentions between the Control and Active Control conditions (Headline Veracity $\times$ Intervention: $b=0.02$, 95% CI [0.02, 0.06][-0.02,\,0.06], $F(1,4046)=1.17$, $p=.28$), and both treatments significantly increased sharing discernment relative to the controls (Headline Veracity $\times$ Intervention: Treatment, $b=0.14$, 95% CI [0.08, 0.20][0.08,\,0.20], $F(1,7208)=23.30$, $p<.001$; Importance Treatment, $b=0.12$, 95% CI [0.07, 0.17][0.07,\,0.17], $F(1,7208)=21.80$, $p<.001$). See Table 22 for the full regression results.
Figure 22: Headline Veracity, Intervention and Participant Type on Twins’ Sharing Intentions.
@@
These additional analyses strengthen confidence in the findings and provide a straightforward comparison for humans and their digital twins, separately, with the reported findings from [43].
@@
These findings extend prior research studying the effectiveness of accuracy nudges in combating the spread of misinformation [43]. Critically, accuracy nudges may have limited effectiveness for stopping the spread of false entertainment news. While prior research has shown that nudges can redirect attention toward accuracy and reduce the spread of false information in political contexts [43], this finding does not seem to generalize to more playful, emotionally engaging content. Specifically, even when primed to consider accuracy or reminded of its importance, humans showed no statistically reliable reduction in their willingness to share false versus true entertainment headlines. Even more, when first prompted to rate the entertainingness of unrelated headlines (i.e., Active Control condition), participants were marginally more likely to share false versus true entertainment headlines, indicating worse discernment. Interestingly, in contrast to humans, digital twins responded as the prior work would predict, demonstrating robust improvements in truth discernment in both the Treatment and Importance Treatment conditions.
One possible reason for this divergence is that digital twins, unlike humans, do not experience the hedonic appeal of entertaining content. For humans, entertainment may override epistemic goals—especially when sharing functions as a form of social play or bonding [34]. In contrast, digital twins may more rigidly apply the prioritization embedded in their programming or training data, defaulting to accuracy in the absence of competing motivations. In this way, digital twins act as idealized rational agents—models of how people ought to behave if their sole goal was to share accurate information. But humans operate in a messier reality, where attention is easily hijacked by humor, surprise, or emotional resonance. This gap underscores a broader challenge in misinformation interventions: even the most well-intentioned nudges may fall short when the content itself draws attention away from accuracy.
From a practical standpoint, the results highlight the limits of one-size-fits-all interventions. If humans are less responsive to accuracy nudges in hedonic contexts, future interventions may need to do more than redirect attention—they may need to reshape the perceived function of sharing itself. For example, interventions could incentivize or socially reward accurate sharing [34], rebalancing the tradeoff between entertainment and truth. More ambitiously, platforms might redesign their interfaces to elevate epistemic cues without dampening the joy of social engagement. At a theoretical level, these findings invite deeper reflection on how models of decision-making should incorporate context-dependent goals, and how digital twins can help surface not just what people do, but why they do it.
@@
Participants first responded to a condition-specific prompt for two minutes. For the gratitude domain, participants received either a gratitude prime (“Please recall carefully and in detail a specific experience in the past when you felt sincerely grateful for someone’s kindness or help”) or baseline prompt (“Please recall carefully and in detail the sequence of your morning routine, such as brushing teeth or changing clothes”), using procedures from [17, 16, 39]. In the lack of control domain, participants received either the lack of control prime (“Please recall a particular incident in which something happened and you did not have control over the situation. Please describe the situation in which you felt a lack of control – what happened, how you felt, etc.”) or a baseline prompt (“Please recall a particular incident in which something happened and you were in control of the situation. Please describe the situation in which you felt in control – what happened, how you felt, etc.”), using procedures from [9, 13, 36, 61].
After the writing task, participants completed a manipulation check, followed by proximal dependent measures (closely related to the target affective state) and distal dependent measures (more distantly related to the affective state). For the gratitude domain, the manipulation check assessed gratitude-related emotions (grateful, thankful, appreciative) on a 7-point scale from “not at all” to “extremely.” The proximal measure used the Elevation scale [51, 59], which asked participants to rate their current feelings on six items such as “Feeling optimistic about humanity” using the same 7-point scale. The distal measure employed an adapted version of the Empathic Concern scale [40], measuring agreement with two statements, “I am very concerned about those most vulnerable to the effects of tariffs” and “I feel compassion for those most affected by rising prices or job losses due to tariffs,” on a 7-point scale from “completely disagree” to “completely agree.”
For the lack of control domain, the manipulation check used items from [22], asking participants to rate their agreement with three statements (“I feel in control of my life”, “I am free to live my life how I wish”, “My experiences in life are due to my own actions”, on a 7-point scale from “not at all” to “extremely.” These items were reverse coded for the analysis, but we present the raw means before reverse coding in the graph. The proximal measure was the Fatigue scale [52, 9], consisting of nine items such as “I had a hard time thinking about the event” rated on the same 7-point scale. The distal measure employed the Desire for Predictability scale [60, 36], containing eight items like “would not like to go into a situation without knowing what I can expect from it” rated on a 7-point scale from “completely disagree” to “completely agree.”
@@
Here, we examined the language used when responding to the prompt. Specifically, we examined the language similarity between a human participant’s response and their twin versus a human participant’s response and a randomly selected twin using the BERT metric, defined as “the sum of cosine similarities between their tokens’ embeddings” [64]. In a random-effects regression with language matching as the dependent variable; same-twin versus different-twin pairing as the independent variable; condition as a covariate; and human participant identifier as a random effect, we found that the language between a human participant and their twin was slightly more similar ($M=0.11$, $SD=0.09$) than between a human participant and a random twin ($M=0.13$, $SD=0.10$; $b=0.014$, $SE=0.002$; $t(999)=5.54$; $p<.001$). Figure 27 shows this by condition.
@@
We explored the predictive validity of the minimalism scale in a sample of human respondents versus their digital twins. As in the original paper on consumer minimalism [63], we expect that human respondents and their digital twins who score high on the Minimalist Consumer Scale will prefer minimalist home environments over non-minimalist interiors.
@@
We recruited 200 human respondents on Prolific and their respective twins. In random order, participants completed the 12 randomized items of the Minimalist Consumer Scale ($M_{\text{human}}=4.84$, $SD_{\text{human}}=1.17$, $\alpha_{\text{human}}=.93$; $M_{\text{twin}}=4.52$, $SD_{\text{twin}}=1.47$, $\alpha_{\text{twin}}=.94$), and indicated their preference for four sets of minimalist versus non-minimalist interiors, presented in random order. All stimuli were pretested and validated in Study 7e of the Consumer Minimalism paper [63]).
@@
Context effects are of great interest to marketers and policy makers alike because they suggest that product assortment can change choice. This study focuses on two classic context effects. The attraction effect occurs when a third option - an asymmetrically dominated decoy - is added to a binary choice set, increasing preference for the option that dominates the decoy [25]. The compromise effect occurs when a different type of decoy is added to a binary choice set, one that makes one of the original options appear as a compromise between the other two, increasing preference for the compromise option [54].
@@
This study focuses on two default effects: the organ donation paradigm by [31], and a green energy default paradigm promoting sustainable energy choices [48]. In the organ donation study, participants who were defaulted to be donors were significantly more likely to remain enrolled than those who had to actively opt in. Similarly, in the green energy study, defaulting participants into using green energy suppliers increased adoption rates compared to requiring them to opt in.
@@
Notably, the organ donation task used the original stimuli from [31], while the green energy default task was slightly modified by changing the company names for the affordable and sustainable options.
@@
Large language models (LLMs) like ChatGPT can accurately describe default effects and reference the relevant literature, presumably because these concepts are well represented in their training data. Default effects, while often robust, vary considerably in magnitude [30]. In the present study, human participants showed a default effect in the green energy adoption paradigm but not in the organ donation paradigm. Digital twins predicted a strong default effect in the organ donation paradigm but no effect in the green energy paradigm.
@@
Luxury brands face the perpetual problem of balancing exclusivity with growth [32]. Diffusion of products into the marketplace inevitably erodes exclusivity and negatively impacts product and brand value [2]. Recent work has demonstrated that a new technology—digital certification—reinforces value when luxury products become diffused [42]. In this study, we aim to replicate this pattern with human participants, and test whether their digital twins respond similarly to digital certification.
Specifically, we test whether attaching a digital passport to a luxury product that has become diffused increases its perceived value (i.e., status perceptions) and monetary value (i.e., price perceptions). Digital passports are a novel Blockchain-based certification system inspired by a recent European Union regulation to create more transparent and sustainable supply chains for physical products [26]. For example, the Aura Blockchain Collective (a non-profit collective of major luxury firms) provides digital passports that allow consumers to follow their purchases through the entire product journey, from the origin of the raw materials to recycling (“Solutions - Aura Blockchain Consortium” n.d.).
@@
We opened the study to 600 human participants on Prolific (46.50% female, 53.50% male; $M_{\text{age}}=49.86$, $SD=14.32$). The study was subsequently run on their 600 digital twins, and the final sample size was 1,200 for our analysis. Participants and their corresponding digital twins were randomly assigned to the same condition in a two-cell (Digital Certification: Digital Passport vs. Control) between-subjects design.
All participants were asked to select one of six luxury sweaters: “Here is a selection of 6 sweaters by a luxury brand. Please choose the sweater that you are most interested in.” Participants were given a description of each sweater, rather than an image, to enable comparison between humans and their digital counterparts, who cannot process visual images (Figure 32, top). The description of each sweater was based on real sweaters designed by Ralph Lauren ($M_{\text{price}}=\mathdollar 531.90$; Figure 32, bottom). To avoid potential confounds, the brand was not mentioned. These six target sweaters were chosen from a set of 20 after an extensive pretest to ensure the products were equally liked, perceived as high-end, suitable for the scenario (i.e., as an outfit for going out), and conspicuous. Additional exploratory analyses controlled for participants’ selected sweater to account for differences across product stimuli (e.g., sweater color, style) that could influence value perceptions.
@@
Participants then responded to the dependent variables by reporting their perceptions of product value (“How expensive do you think the sweater is? (Please enter number of \$)”) and status perceptions (“Please answer the following questions. To what extent do you think the sweater is…”; luxurious, high status, prestigious; 1 = Not at all, 7 = To a great extent; Ward & Dahl, 2014). These three items were averaged for analysis ($\alpha=.94$). We log-normalized the price perception measure to address skewness ($\mu=4.06>2$; i.e., right-skewed; [14]). Finally, participants reported the extent to which they were familiar with digital passports (“How familiar are you with digital passports?”; 1 = Very Unfamiliar, 7 = Very Familiar), the perceived difficulty of the task to account for different prompt lengths across conditions (“How difficult was it to imagine the sweater scenario?”; 1 = Extremely Easy, 7 = Extremely Difficult), and whether they owned real luxury products (0 = No, 1 = Yes).

### v4

@@ Appendix H Centaur Experiments
For humans, accuracy nudges did not improve truth discernment when sharing entertaining news (Figure 21, Panel A). As in (?), there were no significant differences in sharing intentions between the Control and Active Control conditions (Headline Veracity $\times$ Intervention: $b=0.14$, 95% CI [0.07, 0.34][-0.07,\,0.34], $t(2997)=1.31$, $p=.56$). In contrast to the previous findings, neither accuracy nudge significantly increased sharing discernment relative to the controls (e.g., Headline Veracity $\times$ Intervention: Treatment vs. Control, b=0.05b=-0.05, 95% CI [0.25, 0.16][-0.25,\,0.16], t(2997)=0.44t(2997)=-0.44, $p=.97$; Importance Treatment vs. Control, b=0.03b=-0.03, 95% CI [0.23, 0.17][-0.23,\,0.17], t(2997)=0.30t(2997)=-0.30, $p=.99$). See Table LABEL:SI_accuracy_tab3 for pairwise comparisons.
@@
For digital twins, however, accuracy nudges successfully improved truth discernment when sharing entertaining news (Figure 21, Panel B). As in (?), there were no significant differences in sharing intentions between the Control and Active Control conditions (Headline Veracity $\times$ Intervention: b=0.06b=-0.06, 95% CI [0.26, 0.14][-0.26,\,0.14], t(2997)=0.58t(2997)=-0.58, $p=.94$), and both treatments significantly increased sharing discernment relative to the controls (e.g., Headline Veracity $\times$ Intervention: Treatment vs. Control, b=0.61b=-0.61, 95% CI [0.82,0.41][-0.82,\,-0.41], t(2997)=5.93t(2997)=-5.93, $p<.001$; Importance Treatment vs. Control, b=0.44b=-0.44, 95% CI [0.64,0.24][-0.64,\,-0.24], t(2997)=4.27t(2997)=-4.27, $p<.001$). See Table LABEL:SI_accuracy_tab4 for pairwise comparisons.
@@
Table 7: Model B: Random Intercepts and Random Slopes. Headline veracity, intervention, and participant type on sharing. * p<<0.05, ** p<<0.01, *** p<<0.001.
@@
As an additional test, (?) analysis plan was next replicated for humans and their digital twins, separately. While the pre-registered analysis demonstrates sharing intentions for the full range of responses, (?) rescale the continuous sharing intentions into a binary variable capturing responses at or above the midpoint at “likely to share” (= 1) and below the midpoint as “unlikely to share” (= 0), and filtered out participants who would not consider sharing these articles on social media. Because the human dataset revealed significant differences between the Control and Active Control conditions, these groups were not pooled to maintain comparability. This results in a slight deviation from the analysis in (?). Using this framework, their exact analysis was replicated for humans and their digital twins separately, treating the rescaled likelihood to share variable as the DV with Headline Veracity condition, Intervention condition, and their interaction as predictors, applying robust standard errors. Figure 22 shows proportion of sharing across all conditions for humans (Panel A) and their digital twins (Panel B).
@@
Figure 22: Proportion (%) Likely to Share by Headline Veracity and Intervention. NOTE: Shown here is the fraction of “likely’ responses (responses above the midpoint of the six-point Likert scale) by Headline Veracity and Intervention condition; the full distributions of responses are shown below, in figure LABEL:SI_accuracy_3. As with (?), these analyses focus only on participants who indicated that they would consider sharing this social media content. Analysis including all participants does not change the results. Error bars indicate 95% confidence intervals based on robust standard errors clustered on participant and headline.
@@
Replicating the main findings for humans (Model A), accuracy nudges did not improve truth discernment when sharing entertaining news (Figure 22, Panel A). In contrast to (?), sharing discernment was marginally higher in the Active Control compared to the Control condition (Headline Veracity $\times$ Intervention: b=0.03b=-0.03, 95% CI [0.07, 0.00][-0.07,\,0.00], $F(1,5132)=3.53$, $p=.06$), and neither treatment significantly increased sharing discernment relative to the controls (Headline Veracity $\times$ Intervention: Treatment, $b=0.01$, 95% CI [0.05, 0.06][-0.05,\,0.06], $F(1,9656)=0.05$, $p=.82$; Importance Treatment, $b=0.01$, 95% CI [0.04, 0.05][-0.04,\,0.05], $F(1,9656)=0.06$, $p=.80$). See Table LABEL:SI_accuracy_3 for the full regression results.
Table 9: Headline Veracity, Intervention and Participant Type on Humans’ Sharing Intentions. * p<<0.05, ** p<<0.01, *** p<<0.001. When reporting the effects, separate F-tests were conducted; therefore, the reported F statistics are not included in this table, following the reporting approach used by (?).
Humans: Participants who are willing to share content
Model 1: Controls only
Predictors Estimate SE 95% CI t
(Intercept) 0.45*** 0.03 [0.39, 0.50] 17.82
Active Control 0.04 0.02 [0.01, 0.09] 1.64
Headline Veracity -0.01 0.04 [-0.10, 0.08] -0.26
Active Control $\times$ Headline Veracity -0.03 0.02 [-0.07, 0.00] -1.88
Observations 5,136
$R^{2}$ / Adjusted $R^{2}$ 0.003 / 0.003
Model 2: All conditions
Predictors Estimate SE 95% CI t
(Intercept) 0.45*** 0.03 [0.39, 0.50] 17.82
Active Control 0.04 0.02 [-0.01, 0.09] 1.64
Headline Veracity -0.01 0.04 [-0.10, 0.08] -0.26
Active Control $\times$ Headline Veracity -0.03 0.02 [-0.07, 0.00] -1.88
Treatment -0.01 0.03 [-0.07, 0.06] -0.22
Importance Treatment -0.06* 0.03 [-0.12, 0.00] -2.19
Treatment $\times$ Headline Veracity 0.01 0.03 [-0.05, 0.07] 0.23
Importance Treatment $\times$ Headline Veracity 0.01 0.02 [-0.04, 0.06] 0.25
Observations 9,664
$R^{2}$ / Adjusted $R^{2}$ 0.009 / 0.008
@@
Again, replicating the main findings for digital twins (Model A), accuracy nudges improved truth discernment when sharing entertaining news (Figure 22, Panel B). As in (?), there were no significant differences in sharing intentions between the Control and Active Control conditions (Headline Veracity $\times$ Intervention: $b=0.02$, 95% CI [0.02, 0.06][-0.02,\,0.06], $F(1,4046)=1.17$, $p=.28$), and both treatments significantly increased sharing discernment relative to the controls (Headline Veracity $\times$ Intervention: Treatment, $b=0.14$, 95% CI [0.08, 0.20][0.08,\,0.20], $F(1,7208)=23.30$, $p<.001$; Importance Treatment, $b=0.12$, 95% CI [0.07, 0.17][0.07,\,0.17], $F(1,7208)=21.80$, $p<.001$). See Table LABEL:SI_accuracy_4 for the full regression results.
Table 10: Headline Veracity, Intervention and Participant Type on Twins’ Sharing Intentions. * p<<0.05, ** p<<0.01, *** p<<0.001. When reporting the effects, separate F-tests were conducted; therefore, the reported F statistics are not included in this table, following the reporting approach used by (?).
Twins: Participants who are willing to share content
Model 3: Controls only
Predictors Estimate SE CI t
(Intercept) 0.53*** 0.03 [0.46, 0.60] 15.97
Active Control 0.04** 0.01 [0.01, 0.07] 3.05
Headline Veracity -0.01 0.08 [-0.18, 0.15] -0.20
Active Control $\times$ Headline Veracity 0.02 0.02 [0.02, 0.07] 1.08
Observations 4,064
$R^{2}$ / Adjusted $R^{2}$ 0.018 / 0.017
Model 4: All conditions
Predictors Estimate SE CI t
(Intercept) 0.53*** 0.03 [0.46, 0.60] 15.97
Active Control 0.04** 0.01 [0.01, 0.07] 3.05
Headline Veracity -0.01 0.08 [-0.18, 0.15] -0.20
Active Control $\times$ Headline Veracity 0.02 0.02 [0.02, 0.07] 1.08
Treatment -0.07* 0.03 [-0.13, -0.02] -2.81
Importance Treatment 0.10** 0.02 [0.05, 0.15] -4.07
Condition $\times$ Headline Veracity 0.14*** 0.03 [0.08, 0.20] 4.83
Importance Treatment $\times$ Headline Veracity 0.12*** 0.03 [0.06, 0.17] 4.67
Observations 7,216
$R^{2}$ / Adjusted $R^{2}$ 0.058 / 0.057
@@
These additional analyses strengthen confidence in the findings and provide a straightforward comparison for humans and their digital twins, separately, with the reported findings from (?).
@@
These findings extend prior research studying the effectiveness of accuracy nudges in combating the spread of misinformation (?). Critically, accuracy nudges may have limited effectiveness for stopping the spread of false entertainment news. While prior research has shown that nudges can redirect attention toward accuracy and reduce the spread of false information in political contexts (?), this finding does not seem to generalize to more playful, emotionally engaging content. Specifically, even when primed to consider accuracy or reminded of its importance, humans showed no statistically reliable reduction in their willingness to share false versus true entertainment headlines. Even more, when first prompted to rate the entertainingness of unrelated headlines (i.e., Active Control condition), participants were marginally more likely to share false versus true entertainment headlines, indicating worse discernment. Interestingly, in contrast to humans, digital twins responded as the prior work would predict, demonstrating robust improvements in truth discernment in both the Treatment and Importance Treatment conditions.
One possible reason for this divergence is that digital twins, unlike humans, do not experience the hedonic appeal of entertaining content. For humans, entertainment may override epistemic goals, especially when sharing functions as a form of social play or bonding (?). In contrast, digital twins may more rigidly apply the prioritization embedded in their programming or training data, defaulting to accuracy in the absence of competing motivations. In this way, digital twins act as idealized rational agents: models of how people ought to behave if their sole goal was to share accurate information. In the same vein, twins seem more prone to prioritizing “should” or socially desirable characteristics (e.g., source trustworthiness) over truly social characteristics (e.g., entertainingness). But humans operate in a messier reality, where attention is easily hijacked by humor, surprise, or emotional resonance. This gap underscores a broader challenge in misinformation interventions: even the most well-intentioned nudges may fall short when the content itself draws attention away from accuracy.
From a practical standpoint, the results highlight the limits of one-size-fits-all interventions. If humans are less responsive to accuracy nudges in hedonic contexts, future interventions may need to do more than redirect attention—they may need to reshape the perceived function of sharing itself. For example, interventions could incentivize or socially reward accurate sharing (?), rebalancing the tradeoff between entertainment and truth. More ambitiously, platforms might redesign their interfaces to elevate epistemic cues without dampening the joy of social engagement. At a theoretical level, these findings invite deeper reflection on how models of decision-making should incorporate context-dependent goals, and how digital twins can help surface not just what people do, but why they do it.
@@
Participants first responded to a condition-specific prompt for two minutes. For the gratitude domain, participants received either a gratitude prime (“Please recall carefully and in detail a specific experience in the past when you felt sincerely grateful for someone’s kindness or help”) or baseline prompt (“Please recall carefully and in detail the sequence of your morning routine, such as brushing teeth or changing clothes”), using procedures from (?, ?, ?). In the lack of control domain, participants received either the lack of control prime (“Please recall a particular incident in which something happened and you did not have control over the situation. Please describe the situation in which you felt a lack of control – what happened, how you felt, etc.”) or a baseline prompt (“Please recall a particular incident in which something happened and you were in control of the situation. Please describe the situation in which you felt in control – what happened, how you felt, etc.”), using procedures from (?, ?, ?, ?).
After the writing task, participants completed a manipulation check, followed by proximal dependent measures (closely related to the target affective state) and distal dependent measures (more distantly related to the affective state). For the gratitude domain, the manipulation check assessed gratitude-related emotions (grateful, thankful, appreciative) on a 7-point scale from “not at all” to “extremely.” The proximal measure used the Elevation scale (?, ?), which asked participants to rate their current feelings on six items such as “Feeling optimistic about humanity” using the same 7-point scale. The distal measure employed an adapted version of the Empathic Concern scale (?), measuring agreement with two statements, “I am very concerned about those most vulnerable to the effects of tariffs” and “I feel compassion for those most affected by rising prices or job losses due to tariffs,” on a 7-point scale from “completely disagree” to “completely agree.”
For the lack of control domain, the manipulation check used items from (?), asking participants to rate their agreement with three statements (“I feel in control of my life”, “I am free to live my life how I wish”, “My experiences in life are due to my own actions”, on a 7-point scale from “not at all” to “extremely.” These items were reverse coded for the analysis, but we present the raw means before reverse coding in the graph. The proximal measure was the Fatigue scale (?, ?), consisting of nine items such as “I had a hard time thinking about the event” rated on the same 7-point scale. The distal measure employed the Desire for Predictability scale (?, ?), containing eight items like “would not like to go into a situation without knowing what I can expect from it” rated on a 7-point scale from “completely disagree” to “completely agree.”
@@
Here, we examined the language used when responding to the prompt. Specifically, we examined the language similarity between a human participant’s response and their twin versus a human participant’s response and a randomly selected twin using the BERT metric, defined as “the sum of cosine similarities between their tokens’ embeddings” (?). In a random-effects regression with language matching as the dependent variable; same-twin versus different-twin pairing as the independent variable; condition as a covariate; and human participant identifier as a random effect, we found that the language between a human participant and their twin was slightly more similar ($M=0.11$, $SD=0.09$) than between a human participant and a random twin ($M=0.13$, $SD=0.10$; $b=0.014$, $SE=0.002$; $t(999)=5.54$; $p<.001$). Figure 27 shows this by condition.
@@
We explored the predictive validity of the minimalism scale in a sample of human respondents versus their digital twins. As in the original paper on consumer minimalism (?), we expect that human respondents and their digital twins who score high on the Minimalist Consumer Scale will prefer minimalist home environments over non-minimalist interiors.
@@
We recruited 200 human respondents on Prolific and their respective twins. In random order, participants completed the 12 randomized items of the Minimalist Consumer Scale ($M_{\text{human}}=4.84$, $SD_{\text{human}}=1.17$, $\alpha_{\text{human}}=.93$; $M_{\text{twin}}=4.52$, $SD_{\text{twin}}=1.47$, $\alpha_{\text{twin}}=.94$), and indicated their preference for four sets of minimalist versus non-minimalist interiors, presented in random order. All stimuli were pretested and validated in Study 7e of the Consumer Minimalism paper (?)).
@@
Context effects are of great interest to marketers and policy makers alike because they suggest that product assortment can change choice. This study focuses on two classic context effects. The attraction effect occurs when a third option - an asymmetrically dominated decoy - is added to a binary choice set, increasing preference for the option that dominates the decoy (?). The compromise effect occurs when a different type of decoy is added to a binary choice set, one that makes one of the original options appear as a compromise between the other two, increasing preference for the compromise option (?).
@@
This study focuses on two default effects: the organ donation paradigm by (?), and a green energy default paradigm promoting sustainable energy choices (?). In the organ donation study, participants who were defaulted to be donors were significantly more likely to remain enrolled than those who had to actively opt in. Similarly, in the green energy study, defaulting participants into using green energy suppliers increased adoption rates compared to requiring them to opt in.
@@
Notably, the organ donation task used the original stimuli from (?), while the green energy default task was slightly modified by changing the company names for the affordable and sustainable options.
@@
Large language models (LLMs) like ChatGPT can accurately describe default effects and reference the relevant literature, presumably because these concepts are well represented in their training data. Default effects, while often robust, vary considerably in magnitude (?). In the present study, human participants showed a default effect in the green energy adoption paradigm but not in the organ donation paradigm. Digital twins predicted a strong default effect in the organ donation paradigm but no effect in the green energy paradigm.
@@
Luxury brands face the perpetual problem of balancing exclusivity with growth (?). Diffusion of products into the marketplace inevitably erodes exclusivity and negatively impacts product and brand value (?). Recent work has demonstrated that a new technology—digital certification—reinforces value when luxury products become diffused (?). In this study, we aim to replicate this pattern with human participants, and test whether their digital twins respond similarly to digital certification.
Specifically, we test whether attaching a digital passport to a luxury product that has become diffused increases its perceived value (i.e., status perceptions) and monetary value (i.e., price perceptions). Digital passports are a novel Blockchain-based certification system inspired by a recent European Union regulation to create more transparent and sustainable supply chains for physical products (?). For example, the Aura Blockchain Collective (a non-profit collective of major luxury firms) provides digital passports that allow consumers to follow their purchases through the entire product journey, from the origin of the raw materials to recycling.181818“Solutions - Aura Blockchain Consortium” (n.d.), (accessed February 18, 2025), [available at https://auraconsortium.com/solutions].
@@
We opened the study to 600 human participants on Prolific (46.50% female, 53.50% male; $M_{\text{age}}=49.86$, $SD=14.32$). The study was subsequently run on their 600 digital twins, and the final sample size was 1,200 for our analysis. Participants and their corresponding digital twins were randomly assigned to one of two conditions in a two-cell (Digital Certification: Digital Passport vs. Control) between-subjects design (each human and their digital twin were assigned to the same condition).
All participants were asked to select one of six luxury sweaters: “Here is a selection of 6 sweaters by a luxury brand. Please choose the sweater that you are most interested in.” Digital twins were given a textual description of each sweater, rather than an image, to enable comparison between humans and their digital counterparts, who cannot process visual images (Figure 32, top). The description of each sweater was based on real sweaters designed by Ralph Lauren ($M_{\text{price}}=\mathdollar 531.90$; Figure 32, bottom). To avoid potential confounds, the brand was not mentioned. These six target sweaters were chosen from a set of 20 after an extensive pretest to ensure the products were equally liked, perceived as high-end, suitable for the scenario (i.e., as an outfit for going out), and conspicuous. Additional exploratory analyses controlled for participants’ selected sweater to account for differences across product stimuli (e.g., sweater color, style) that could influence value perceptions.
@@
Participants then responded to the dependent variables by reporting their perceptions of product value (“How expensive do you think the sweater is? (Please enter number of \$)”) and status perceptions (“Please answer the following questions. To what extent do you think the sweater is…”; luxurious, high status, prestigious; 1 = Not at all, 7 = To a great extent; Ward & Dahl, 2014). These three items were averaged for analysis ($\alpha=.94$). We log-normalized the price perception measure to address skewness ($\mu=4.06>2$; i.e., right-skewed; (?)). Finally, participants reported the extent to which they were familiar with digital passports (“How familiar are you with digital passports?”; 1 = Very Unfamiliar, 7 = Very Familiar), the perceived difficulty of the task to account for different prompt lengths across conditions (“How difficult was it to imagine the sweater scenario?”; 1 = Extremely Easy, 7 = Extremely Difficult), and whether they owned real luxury products (0 = No, 1 = Yes).